If the charge for a taxi ride is $2.25 for the first 1/2 mile and $0.75 for each additional 1/8 mile, how many miles did the taxi travel for a ride that cost $12.00?
Answer by lwsshak3(6480)

If the charge for a taxi ride is $2.25 for the first 1/2 mile and $0.75 for each additional 1/8 mile, how many miles did the taxi travel for a ride that cost $12.00?
...
let x=miles traveled
traveled 1/2 mile for $2.25
traveled x-1/2 miles at a cost of $.75 for each 1/8 mile
..
2.25+.75(x-1/2)/(1/8)=12
2.25+.75(x-1/2)*8=12
2.25+6(x-1/2)=12
2.25+6x-3=12
6x-.75=12
6x=12.75
x=12.75/6=2.125 or 2+1/8 miles (ans)